Calculate the partition function, free energy, entropy, CV and Cp of N linear 

harmonic oscillators


1
Expert's answer
2021-04-06T06:33:04-0400

We know that Maxwell distribution function

ni= gi"e^{-\alpha }e^{-\beta\epsilon}"

ni=gi"/e^{(\alpha+\beta\epsilon)}"

Total no of partical

N="\Sigma" ni

N="\Sigma" gi"e^{-\alpha}e^{-\beta\epsilon}"

N=A"\Sigma"gi"e^{-\beta\epsilon}"

"\frac{N}{A}=\Sigma" gi"e^{\Sigma-\beta\epsilon}" Where gi degenracy

Z="\Sigma e^{-\beta\epsilon}"

Where n D harmonic oscillator energy

"\epsilon" =E="(n+1\/2)"


Free energy

F= -KB"TlnZ"

F=-KBT"\ln\frac{k(B)T}{hw}"

Entropy

S=KB"\frac{d}{dT}(T\ln Z)"

Cp=Cv="\frac{d}{dT}(\overline E)" =KB

"\overline E =KT^2\frac{d}{dT}\ln Z"